# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

### v2

@@ Abstract
An integrated sensing and communication (ISAC) design for pinching antenna systems (PASS) is proposed, where the pinching antennas are deployed for establishing reliable line-of-sight communication and sensing links. More particularly, a separated ISAC design is proposed for the two-waveguide PASS, where one waveguide is used to emit the joint communication and sensing signals while the other waveguide is used to receive the reflected echo signals. Based on this framework, a penalty-based alternating optimization algorithm is proposed to maximize the illumination power as well as ensure the communication quality-of-service requirement. Numerical results demonstrate that 1) the proposed PASS-ISAC scheme outperforms the other baseline schemes, and 2) the considered equal power allocation model achieves an upper bound performance.
@@ I Introduction
Fuelled by the burgeoning demands for massive data transmission and pervasive network coverage, flexible antennas have emerged as a promising technique for sixth-generation (6G) cellular systems. Benefiting from their ability to reconfigure the wireless channel, flexible antennas can significantly enhance the throughput of wireless networks. However, traditional flexible antennas (e.g., moveable antennas [1] and fluid antennas [2]) merely permit the adjustment of the antenna position within a range of orders of magnitude comparable to the carrier wavelength. Against this backdrop, the pinching antenna has emerged [3], which is a type of dielectric waveguide-based leaky wave antenna. By applying dielectric particles to a particular point on the dielectric waveguide, a pinching antenna can be activated to establish EM radiation fields and form a communication area [4]. Then, the EM signal inside the dielectric waveguide will be radiated from the pinching antenna to free space with a defined phase shift adjustment (referred to as the pinching beamformer). Notably, as the dielectric waveguide can be pinched at any position to radiate radio waves, the pinching antenna can flexibly move along the dielectric waveguide over a length of dozens of meters, thereby relocating to the closest position to the receiver and creating reliable LoS links.
@@
To enable emerging applications, such as autonomous driving, extended reality, and Metaverse, sensing functionality is recognized as an important indicator of future networks. In pursuit of this vision, the integrated sensing and communication (ISAC) technology has drawn significant attention recently [5], which aims to leverage the cellular network hardware platforms and dedicated signal processing algorithms to achieve the incorporation of communication and sensing functionalities. Recently, it has been claimed that conducting ISAC transmission in the pinching antenna systems (PASS) can further upgrade the communication and sensing (C&S) performance of the network [6]. On the one hand, the pinching antenna can be flexibly repositioned to augment the echo signal energy. On the other hand, the wide-range mobility characteristic of pinching antennas results in an antenna aperture spanning dozens of meters. It inherently enables near-field sensing, e.g., the possibility of simultaneous angular and distance information estimation and even target velocity sensing, thereby offering a more comprehensive and accurate sensing of the surrounding environment. Nevertheless, as of the present moment, research in the PASS-ISAC remains conspicuously absent.
@@
Motivated by the above, this paper proposes a separated ISAC design for PASS. To elaborate, the base station (BS) is connected with two dielectric waveguides, where one waveguide is used to transmit the communication and sensing signals, while the other is employed to collect the reflected echo signals from the target. We aim to maximize the illumination power at the target while satisfying the quality-of-service (QoS) requirement of the communication user by optimizing the pinching beamforming offered by the mobility of pinching antennas. A penalty-based alternating optimization (AO) algorithm is proposed to handle the non-convex optimization problem, where the optimal radiation pattern is obtained under an arbitrary antenna activation configuration. Numerical results evaluate the superiority of the proposed scheme over the baseline schemes. It also reveals that the equal power model achieves a performance comparable to the upper bound.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a PASS-ISAC system, where a dual-function BS conveys with a single-antenna communicating user while sensing a point-like target. The BS is connected with two dielectric waveguides, where one waveguide is used to emit the downlink signals, and the other is employed to receive the reflected echo signals from the target. It is assumed that each dielectric waveguide consists of $N$ pinching antennas. We aim to achieve simultaneous C&S transmission in PASS using only information-bearing signals, which are emitted from the transmitting antennas. Then, the reflected echoes from the target would be collected at the receiving antennas, which are transmitted to the BS for parameter estimation.
@@
A three-dimensional (3D) coordination system is considered, where two dielectric waveguides extended from the BS are assumed to be parallel to the x-axis with respect to the x-o-y plane at a distance $d$. The position of the $n$th pinching antenna distributed along the transmitting and receiving dielectric waveguides can be denoted as $\psi_{n}^{\text{p}}=(x_{n}^{\text{p}},0,d)$ and $\psi_{n}^{\text{q}}=(x_{n}^{\text{q}},\bar{y},d)$. The communication user and sensing target are located in the x-o-y plane, with the coordinates of $\psi^{\text{c}}=(x_{1},y_{1},0)$ and $\psi^{\text{s}}=(x_{2},y_{2},0)$, respectively. Furthermore, we assume the target is a static node or moves at a low speed. Thus, the Doppler effect is neglected in this work.
@@
In the considered network, the pinching antennas are non-uniformly disposed on the dielectric waveguide covering the entire range of the user’s activity, which implies that the aperture of pinching antennas may have the same order of magnitude as the signal transmission distance. Without loss of accuracy, we adopt the spherical-wave-based near-field channel model, where only the LoS path is considered. Let $r_{\text{s}}$ and $\phi_{\text{s}}$ denote the distance of the target relative to the origin of the coordinate system and the azimuth, respectively. The coordinate of the target can be rewritten as $(r_{\text{s}}\cos\phi_{\text{s}},r_{\text{s}}\sin\phi_{\text{s}},0)$. Consequently, the distance from the $n$-th pinching antenna to the target is given by
$r_{n}(r_{\text{s}},\phi_{\text{s}})$ $=\|\psi^{\text{s}}-\psi_{n}^{\text{p}}\|$
$=\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}.$ (1)
@@
Thus, the channel vector from the transmitting antenna arrays to the target and the communication user can be expressed as
@@
$\mathbf{h}_{\text{s}}(r_{\text{s}},\phi_{\text{s}})=\left[\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}r_{1}(r_{\text{s}},\phi_{\text{s}})}}{r_{1}(r_{\text{s}},\phi_{\text{s}})},\cdots,\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}r_{N}(r_{\text{s}},\phi_{\text{s}})}}{r_{N}(r_{\text{s}},\phi_{\text{s}})}\right]^{T},$ (2)
@@
$\mathbf{h}_{\text{c}}=\left[\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\|\psi^{\text{c}}-\psi_{1}^{\text{p}}\|}}{\|\psi^{\text{c}}-\psi_{1}^{\text{p}}\|},\cdots,\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\|\psi^{\text{c}}-\psi_{N}^{\text{p}}\|}}{\|\psi^{\text{c}}-\psi_{N}^{\text{p}}\|}\right]^{T},$ (3)
@@
where $\lambda=\frac{2\pi}{f_{\text{c}}}$ dentes the wavelength, $f_{\text{c}}$ is the frequency of the carrier wave, $\eta=\frac{c^{2}}{16\pi^{2}f_{\text{c}}^{2}}$, and $c$ denotes the speed of light.
II-A ISAC Model for Pinching Antenna Networks
@@
Consider a coherent time block of length $T$, the BS transmits the superimposed communication and sensing waveforms to the dielectric waveguide. We assume the communication channel condition and the sensing parameters remain unchanged during one coherent time block. Thus, the emitted signal at the $t$-th time slot is given by $s(t)\in\mathbb{C}$, which is assumed to be normalized and independently distributed, i.e., $\mathbb{E}\{|s(t)|^{2}\}=1$ and $\mathbb{E}\{s(t)s^{*}(\bar{t})\}=0$. On receiving $s(t)$, the dielectric waveguide radiates the signal as follows.
@@
$\mathbf{x}(t)=\left[\sqrt{P_{1}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N}}e^{\jmath\theta_{N}}\right]s(t),$ (4)
where $P_{n}$ and $\theta_{n}$ denote the power allocation and the radiation phase shift at the $n$-th pinching antenna. For ease of implementation, the equal power allocation model is assumed, i.e., $P_{1}=\cdots=P_{N}=\frac{P_{\text{max}}}{N}$ [4]. $\theta_{n}$ is defined by $2\pi\eta_{\text{eff}}\frac{\|\psi_{0}^{\text{p}}-\psi_{n}^{\text{p}}\|}{\lambda}$, where $\psi_{0}^{\text{p}}$ denotes the location of the feed point, and $\eta_{\text{eff}}$ denotes the effective refractive index of the dielectric waveguide. According to (2)-(4), we can observe that the positions of pinching antennas have a significant impact on both the free-space path loss of wireless channels and the phase shift of radiated signals. This indicates that it is possible to establish the favorable wireless propagation and manipulate the radiated characteristics of signals by altering the positions of pinching antennas. The capability of reconfiguring signal propagation is referred as pinching beamforming.
To facilitate the beamforming design, we consider a semi-continuous activation structure for pinching antennas. As shown in Fig. 2, the pinching antennas can be activated in a large-scale range over the dielectric waveguide to change the free-space path loss and the spherical-wave array response. However, when the large-scale activation position is determined, the pinching antennas can also be adjusted in a small-scale range to change the phase shift of the radiation signal, but without significantly affecting the wireless channel characteristics. For example, consider a millimeter wave ISAC scenario with a carrier frequency of $28$ GHz, a dielectric waveguide length of 50 meters (m), and $\eta_{\text{eff}}=1.4$, the small-scale phase-shift adjustment range will not exceed a maximum of $7.7$ millimeter, which has a negligible impact on the large-scale path loss as well as the spherical-wave array response. Consequently, it can be reasonably assumed that the radiation phase shift design of pinching antennas is independent of the large-scale activation position optimization.
Figure 2: Semi-continuous activation structure.
In this paper, we aim to utilize the communication signal to achieve simultaneous communication and target sensing. Thus, the transmitted signal can be rewritten as $\mathbf{x}(t)=\mathbf{w}s(t)$, where $\mathbf{w}=\left[\sqrt{P_{1}\alpha_{\text{c}}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N}\alpha_{\text{c}}}e^{\jmath\theta_{N}}\right]$.
II-A1 Communication Performance Metric
@@
With the aforementioned signal model, the received signals at the communication user are given by
$y(t)=\mathbf{h}_{\text{c}}^{H}\mathbf{w}s(t)+n(t),$ (5)
@@
where $n(t)\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) at the communication user. Hence, the achievable rate of the communication user is given by
$R=\log_{2}\left(1+\frac{|\mathbf{h}_{\text{c}}^{H}\mathbf{w}|^{2}}{\sigma^{2}}\right).$ (6)
II-A2 Sensing Performance Metric
@@
For target sensing, we adopt the illumination power as the performance metric, which characterizes the received sensing signal power at the target [7]. Thus, the illumination power with respect to azimuth angle $\phi_{\text{s}}$ and distance $r_{\text{s}}$ is given by
$P(\theta_{\text{s}},r_{\text{s}})$ $=\mathbb{E}\left\{\left|\mathbf{h}_{\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{w}s(t)\right|^{2}\right\}$
$=\mathbf{h}_{\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{\text{s}}.$ (7)
II-B Problem Formulation

### v3

@@ Abstract
An integrated sensing and communication (ISAC) design for pinching antenna systems (PASS) is proposed, where the pinching antennas are deployed to establish reliable line-of-sight communication and sensing links. More particularly, a separated ISAC design is proposed for the two-waveguide PASS, where one waveguide is used to emit the information-bearing signals for ISAC transmission while the other waveguide is used to receive the reflected echo signals. Based on this framework, a penalty-based alternating optimization algorithm is proposed to maximize the illumination power as well as ensure the communication quality-of-service requirement. Numerical results demonstrate that the proposed PASS-ISAC scheme outperforms the conventional antenna scheme.
@@ I Introduction
Fuelled by the burgeoning demands for massive data transmission and pervasive network coverage, flexible antennas have emerged as a promising technique for sixth-generation (6G) cellular systems. Benefiting from their ability to reconfigure the wireless channel, flexible antennas can significantly enhance the throughput of wireless networks. However, traditional flexible antennas (e.g., movable antennas [1] and fluid antennas [2]) merely permit the adjustment of the antenna position within a range of orders of magnitude comparable to the carrier wavelength. Against this backdrop, the pinching antenna has emerged [3], which is a type of dielectric waveguide-based leaky wave antenna. By applying dielectric particles to a particular point on the dielectric waveguide, a pinching antenna can be activated to establish EM radiation fields and form a communication area [4]. Then, the EM signal inside the dielectric waveguide will be radiated from the pinching antenna to free space with a defined phase shift adjustment (referred to as the pinching beamformer). Notably, as the dielectric waveguide can be pinched at any position to radiate radio waves, the pinching antenna can flexibly move along the dielectric waveguide over a length of dozens of meters, thereby relocating to the closest position to the receiver and creating reliable LoS links.
@@
To enable emerging applications, such as autonomous driving, extended reality, and the Metaverse, sensing functionality is recognized as an important indicator of future networks. In pursuit of this vision, the integrated sensing and communication (ISAC) technology has drawn significant attention recently [5], which aims to leverage the cellular network hardware platforms and dedicated signal processing algorithms to achieve the incorporation of communication and sensing functionalities. Recently, it has been claimed that conducting ISAC transmission in the pinching antenna systems (PASS) can further upgrade the communication and sensing (C&S) performance of the network [6]. On the one hand, the pinching antenna can be flexibly repositioned to augment the echo signal energy. On the other hand, the wide-range mobility characteristic of pinching antennas results in an antenna aperture spanning dozens of meters. It inherently enables near-field sensing, e.g., the possibility of simultaneous angular and distance information estimation and even target velocity sensing, thereby offering a more comprehensive and accurate sensing of the surrounding environment. Nevertheless, as of the present moment, research in the PASS-ISAC remains conspicuously absent.
@@
Motivated by the above, this paper proposes a separated ISAC design for PASS. To elaborate, the base station (BS) is connected with two dielectric waveguides, where one waveguide is used to transmit the downlink signals, while the other is employed to collect the reflected echo signals from the target. We aim to maximize the illumination power at the target while satisfying the quality-of-service (QoS) requirement of the communication user by optimizing the pinching beamforming offered by the mobility of pinching antennas. A penalty-based alternating optimization (AO) algorithm is proposed to handle the non-convex optimization problem, where the positions of pinching antennas are updated in an element-wise manner. Numerical results evaluate the superiority of the proposed scheme over the baseline schemes.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a PASS-ISAC system, where a dual-function BS conveys with a single-antenna communication user while sensing a point-like target. The BS is connected with two dielectric waveguides of length $L$, each of which consists of $N$ pinching antennas. To achieve the simultaneous C&S transmission, a separated ISAC design is proposed. Specifically, the downlink information-bearing signals are emitted from one waveguide (referred to as transmitting antennas). Then, the reflected echoes from the target would be collected at the other waveguide (referred to as receiving antennas), which are further transmitted to the BS for parameter estimation.
@@
A three-dimensional (3D) coordination system is considered, where two dielectric waveguides extended from the BS are assumed to be parallel to the x-axis with respect to the x-o-y plane at a height $d$. The position of the $n$-th pinching antenna distributed along the transmitting and receiving dielectric waveguides can be denoted as $\psi_{n}^{\text{p}}=(x_{n}^{\text{p}},0,d)$ and $\psi_{n}^{\text{q}}=(x_{n}^{\text{q}},y^{\text{q}},d)$. The communication user and sensing target are located in the x-o-y plane. Let $r_{\text{c}}$ and $\varphi_{\text{c}}$ denote the distance and the azimuth angle of the communication user relative to the origin of the coordinate system. Thus, the coordinates of communication user is given by $\psi^{\text{c}}=(r_{\text{c}}\cos\varphi_{\text{c}},r_{\text{c}}\sin\varphi_{\text{c}},0)$. Similarly, the target is located in $\psi^{\text{s}}=(r_{\text{s}}\cos\varphi_{\text{s}},r_{\text{s}}\sin\varphi_{\text{s}},0)$. Furthermore, we assume the target is a static node or moves at a low speed. Thus, the Doppler effect is neglected in this work.
II-A Channel Model
@@
In the considered network, the pinching antennas are non-uniformly disposed on the dielectric waveguide covering the entire range of the user’s activity, which implies that the aperture of the pinching antennas may have the same order of magnitude as the signal transmission distance. Without loss of accuracy, we adopt the spherical-wave-based near-field channel model, where only the LoS path is considered. Consequently, the distance from the $n$-th pinching antenna to the target is given by
$r_{n}^{\zeta}(r_{\zeta},\varphi_{\zeta})$ $=\|\psi^{\zeta}-\psi_{n}^{\text{p}}\|$
$=\sqrt{r_{\zeta}^{2}-2r_{\zeta}\cos\varphi_{\zeta}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}},\ \zeta\in\{\text{s},\text{c}\},$ (1)
@@
Thus, the free space channel vector from the transmitting antennas to the target and the communication user can be expressed as
@@
$\mathbf{h}_{\text{s}}(\mathbf{x}^{\text{p}})=\left[\frac{\eta^{\frac{1}{2}}e^{-\jmath\frac{2\pi}{\lambda}r_{1}^{\text{s}}(r_{,}\varphi_{\text{s}})}}{r_{1}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})},\cdots,\frac{\eta^{\frac{1}{2}}e^{-\jmath\frac{2\pi}{\lambda}r_{N}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})}}{r_{N}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})}\right]^{H},$ (2)
@@
$\mathbf{h}_{\text{c}}(\mathbf{x}^{\text{p}})=\left[\frac{\eta^{\frac{1}{2}}e^{-\jmath\frac{2\pi}{\lambda}r_{1}^{\text{c}}(r_{,}\varphi_{\text{c}})}}{r_{1}^{\text{c}}(r_{,}\varphi_{\text{c}})},\cdots,\frac{\eta^{\frac{1}{2}}e^{-\jmath\frac{2\pi}{\lambda}r_{N}^{\text{c}}(r_{,}\varphi_{\text{c}})}}{r_{N}^{\text{c}}(r_{,}\varphi_{\text{c}})}\right]^{H},$ (3)
@@
where $\mathbf{x}^{\text{p}}=[x_{1}^{\text{p}},\cdots,x_{N}^{\text{p}}]$ denotes the coordinates of pinching antennas, $\lambda=\frac{c}{f_{\text{c}}}$ denotes the wavelength, $f_{\text{c}}$ is the frequency of the carrier wave, $\eta=\frac{c^{2}}{16\pi^{2}f_{\text{c}}^{2}}$, and $c$ denotes the speed of light.
@@
In this paper, the BS aims to utilize the communication signal to achieve simultaneous communication and target sensing. Consider a coherent time block of length $T$, the communication channel condition and the sensing parameters are assumed to remain unchanged during one coherent time block. Thus, the emitted signal at the $t$-th time slot is given by $s(t)\in\mathbb{C}$, which is assumed to be normalized and independently distributed, i.e., $\mathbb{E}\{|s(t)|^{2}\}=1$ and $\mathbb{E}\{s(t)s^{*}(\bar{t})\}=0$. On receiving $s(t)$, the dielectric waveguide radiates the signal $\mathbf{x}(t)=\sqrt{P_{\text{T}}}\mathbf{g}(\mathbf{x}^{\text{p}})s(t)$, where $\mathbf{g}(\mathbf{x}^{\text{p}})$ denotes the in-waveguide channel and can be expressed as
@@
$\mathbf{g}(\mathbf{x}^{\text{p}})=\left[\sqrt{\alpha_{1}}e^{-\jmath\theta_{1}},\cdots,\sqrt{\alpha_{N}}e^{-\jmath\theta_{N}}\right]^{T},$ (4)
where $\theta_{n}$ denotes the radiation phase shift at the $n$-th pinching antenna, and $P_{\text{T}}$ denotes the transmit power at the BS. $\alpha_{n}$ denotes the power allocation coefficients at the $n$-th pinching antenna, which can be modeled as the equal power allocation model $\sqrt{\alpha_{n}}=\sqrt{\frac{\alpha_{\text{s}}}{N}}$ [4] or the proportional power allocation model $\sqrt{\alpha_{n}}=\delta(\sqrt{1-\delta^{2}})^{n-1}$ [7]. $\delta=\sqrt{1-(1-\alpha_{\text{s}})^{\frac{1}{N}}}$ represents the proportional coefficient, and $\alpha_{\text{s}}=\sum_{n=1}^{N}\alpha_{n}$ denotes the radiation coefficient of pinching antennas. For ease of implementation, the equal power allocation model is considered in this paper. $\theta_{n}$ is defined by $2\pi\eta_{\text{eff}}\frac{\|\psi_{0}^{\text{p}}-\psi_{n}^{\text{p}}\|}{\lambda}$, where $\psi_{0}^{\text{p}}$ denotes the location of the feed point, and $\eta_{\text{eff}}$ denotes the effective refractive index of the dielectric waveguide.
II-B Signal Model
With the above channel model, it is readily observed that the positions of pinching antennas have a significant impact on both the free space channel $\{\mathbf{h}_{\text{s}}(\mathbf{x}^{\text{p}}),\mathbf{h}_{\text{c}}(\mathbf{x}^{\text{p}})\}$ and the in-waveguide channel $\mathbf{g}(\mathbf{x}^{\text{p}})$. As a result, it becomes possible to establish favorable wireless propagation while manipulating the radiated characteristics of signals by altering the positions of pinching antennas in the PASS. To characterize the two aspects of the signal reconfiguration capabilities of pinching antennas, we refer to it as pinching beamforming in this paper. Let $\mathbf{w}(\mathbf{x}^{\text{p}})$ and $\mathbf{v}(\mathbf{x}^{\text{p}})$ denote the pinching beamforming for the communication user and the sensing target, which are also the functions of $\mathbf{x}^{\text{p}}$. $\mathbf{w}(\mathbf{x}^{\text{p}})$ and $\mathbf{v}(\mathbf{x}^{\text{p}})$ are given by
$\mathbf{w}(\mathbf{x}^{\text{p}})\!=\!\!\left[\!\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{1}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})\!+\!\theta_{1})}}{\frac{1}{\sqrt{\alpha_{1}}}r_{1}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})}\!,\!\cdots\!,\!\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{N}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})\!+\!\theta_{N})}}{\frac{1}{\sqrt{\alpha_{N}}}r_{N}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})}\!\right]^{T}\!\!,$ (5)
$\mathbf{v}(\mathbf{x}^{\text{p}})\!=\!\!\left[\!\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{1}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})\!+\!\theta_{1})}}{\frac{1}{\sqrt{\alpha_{1}}}r_{1}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})}\!,\!\cdots\!,\!\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{N}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})\!+\!\theta_{N})}}{\frac{1}{\sqrt{\alpha_{N}}}r_{N}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})}\!\right]^{T}\!\!.$ (6)
In this paper, we consider an ideal activation model of the pinching antenna, i.e., continuous activation. It indicates that the pinching antennas can be activated at any position of the dielectric waveguide. Thus, the positions of pinching antennas satisfy
$\mathbf{x}^{\text{p}}\in\mathcal{X}=\left\{|x_{n}^{\text{p}}-x_{m}^{\text{p}}|\geq\Delta x\ (n\neq m),\ x_{n}^{\text{p}}\in\left[-\frac{L}{2},\frac{L}{2}\right]\right\},$ (7)
where $\Delta x$ represents the minimum antenna space between two adjacent pinching antennas.
@@ II-B Signal Model
1) Communication Performance Metric: With the aforementioned signal model, the received signals at the communication user are given by
$y(t)$ $=\sqrt{P_{\text{T}}}\mathbf{h}_{\text{c}}^{H}(\mathbf{x}^{\text{p}})\mathbf{g}(\mathbf{x}^{\text{p}})s(t)+n(t)$
$=\sqrt{P_{\text{T}}}\bm{\eta}^{H}\mathbf{w}(\mathbf{x}^{\text{p}})s(t)+n(t),$ (8)
@@
where $\bm{\eta}=[\eta^{\frac{1}{2}},\cdots,\eta^{\frac{1}{2}}]^{T}\in\mathbb{C}^{N\times 1}$ is a constant vector, and $n(t)\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) at the communication user. Hence, the achievable rate of the communication user is given by
$R=\log_{2}\left(1+\frac{P_{\text{T}}|\bm{\eta}^{H}\mathbf{w}(\mathbf{x}^{\text{p}})|^{2}}{\sigma^{2}}\right).$ (9)
@@
2) Sensing Performance Metric: For target sensing, we adopt the illumination power as the performance metric, which characterizes the received sensing signal power at the target [8]. Thus, the illumination power with respect to azimuth angle $\varphi_{\text{s}}$ and distance $r_{\text{s}}$ is given by
$P_{\text{s}}$ $=\mathbb{E}\left\{\left|\sqrt{P_{\text{T}}}\mathbf{h}_{\text{s}}^{H}(\mathbf{x}^{\text{p}})\mathbf{g}(\mathbf{x}^{\text{p}})s(t)\right|^{2}\right\}$
$=P_{\text{T}}\bm{\eta}^{H}\mathbf{v}(\mathbf{x}^{\text{p}})\mathbf{v}^{H}(\mathbf{x}^{\text{p}})\bm{\eta}.$ (10)
II-C Problem Formulation